OpenStudy (samsan9):

find (fog)(2) and (f+g)(2) when f(x)=1/xand g(x)=4x+9

myininaya (myininaya):

So we need to find f(g(2)) and we also need to find f(2)+g(2)

myininaya (myininaya):

do you know how to find g(2)?

OpenStudy (samsan9):

do you plug in 2 for the g=4x+9?

myininaya (myininaya):

you plug in x for 2 in the expression named g yes

OpenStudy (samsan9):

so it equals 17?

myininaya (myininaya):

g(2) is 17 that is right

myininaya (myininaya):

so do you know how to find f(g(2)) you just found g(2) so you need to find f(g(2)) now f(g(2))=f(17) I replaced g(2) with 17 now you need to find f(17) f(g(2))=f(17)=?

OpenStudy (samsan9):

so you plug in 17 where the f(x)=1/x right?

myininaya (myininaya):

right

myininaya (myininaya):

so f(g(2))=?

OpenStudy (samsan9):

1/17?

myininaya (myininaya):

right

myininaya (myininaya):

now f(2)+g(2) you already found g(2) earlier you said it was 17 so we have now f(2)+g(2) f(2)+17

myininaya (myininaya):

any clue what to do now?

OpenStudy (samsan9):

\[\frac{ 1 }{ 17 }+\frac{ 17 }{ 1 }i am assuming\]

myininaya (myininaya):

how did you get f(2) is 1/17? I think f(x)=1/x so f(2)=?

OpenStudy (samsan9):

oh my apoligize so is it 1/2

myininaya (myininaya):

yep great job now you need to simplify f(2)+g(2) f(2)+g(2) = 1/2+17

myininaya (myininaya):

guess you could write that as a mix fraction

myininaya (myininaya):

which basically it is already written that way

myininaya (myininaya):

anyways lets summarize what we did we found f(g(2))=1/17 we also found f(2)+g(2)=17+1/2

OpenStudy (samsan9):

so do we just plug in 2 where ever there is an x and then we just just combine f and g in (fog) or do we just leave it as (fog)(2) and do we just leave it as 1/2+17 for (f+g)(2)?

myininaya (myininaya):

i didn't write 17+1/2 as a mix fraction i left that part for you we were asked to find (f o g)(2) which is f(g(2)) all you do first is take the g function and replace x's with 2's then take whatever result you get from that in plug into f which I believe we got 1/17 if i remember correctly by the way (g o f)(2) means g(f(2)) so you would start with f first if you seen this (f+g)(2) just means f(2)+g(2)

OpenStudy (samsan9):

thank you :D